# footnotetext: * Equal contribution

### v1

@@ Abstract
We introduce Loki, a compute-efficient framework for co-designing morphologies and control policies that generalize across unseen tasks. Inspired by biological adaptation—where animals quickly adjust to morphological changes—our method overcomes the inefficiencies of traditional evolutionary and quality-diversity algorithms. We propose learning convergent functions: shared control policies trained across clusters of morphologically similar designs in a learned latent space, drastically reducing the training cost per design. Simultaneously, we promote divergent forms by replacing mutation with dynamic local search, enabling broader exploration and preventing premature convergence. The policy reuse allows us to explore 780$\sim$ 780$\times$ 780 × more designs using 78% fewer simulation steps and 40% less compute per design. Local competition paired with a broader search results in a diverse set of high-performing final morphologies. Using the UNIMAL design space and a flat-terrain locomotion task, Loki discovers a rich variety of designs—ranging from quadrupeds to crabs, bipedals, and spinners—far more diverse than those produced by prior work. These morphologies also transfer better to unseen downstream tasks in agility, stability, and manipulation domains (e.g. 2×2$\times$2 × higher reward on bump and push box incline tasks). Overall, our approach produces designs that are both diverse and adaptable, with substantially greater sample efficiency than existing co-design methods.
@@ 1 Introduction
Embodied cognition asserts that intelligence emerges not just from the brain but is shaped by the physical embodiment [1]. This perspective acknowledges that “how a creature behaves is influenced heavily by their body,” which acts as an interface between the controller and the environment [2]. A four legged embodiment gallops while a two legged one simply runs. This adaptation occurs across both evolutionary timescales—as in Darwin’s finches, whose beak shapes evolved to exploit different food sources—and individual lifetimes, through mechanisms like Wolff’s Law, where bones strengthen in response to load [3]. Animals can seamlessly readapt their controllers to accommodate even subtle changes in morphology [4].
Designing—or discovering—optimal embodiments for robots is a difficult combinatorial challenge [5, 6, 7, 8, 9, 10, 11, 12, 4, 13]. Co-design, the joint optimization of morphology and control, is typically framed as a bi-level optimization: an outer loop searches for morphologies (often using evolutionary strategies), while an inner loop trains a control policy for each design to evaluate its fitness (performance on a task). Due to the vastness of the design space, exhaustive search is infeasible [14, 4]. Worse, the fitness landscape is often deceptive [15], causing evolutionary strategies to converge prematurely to local optima. Escaping such traps requires maintaining diversity across the population [16, 17, 18]. Quality-diversity algorithms address this by promoting exploration of high-performing yet diverse designs [19, 20, 21, 22]. However, evaluating each candidate is costly, and diversity comes at the expense of the number of designs evaluated in the inner loop [23]. Unlike animals—who quickly adapt controllers to new bodies—robots still require retraining from scratch for each morphology, as policies transfer poorly across embodiments [4, 24].
@@
Second, we break away from how most algorithms sample new designs using mutation. Mutating high-performing designs leads to limited exploration. Instead, we sample divergent forms by adopting dynamic local search (DLS) [25, 26, 27]. We maintain a dynamic elite pool per cluster to train a shared policy for the cluster. During training, random designs are sampled from a cluster, evaluated by the current policy, and top ones replace the worst-performing elites in the pool (see Alg. 1). Compared to mutation, which typically yields small, incremental changes, DLS enables broader and more adaptive exploration, reducing the risk of getting stuck in local minima. Since evaluations are made inexpensive with shared policies, we can afford to use random search instead of local mutation, efficiently filtering out low-quality candidates.
We use UNIMAL [4], an expressive design space encompassing approximately $10^{18}$ unique morphologies with fewer than 10 limbs. We evolve morphologies on a simple flat-terrain locomotion task. In nature, locomotion acts as a universal selection pressure across species. It is task-agnostic—helping to prevent overfitting to narrow objectives—and is efficient to simulate and easy to reward.
Beyond being substantially more efficient, Loki discovers both genetically and behaviorally diverse high-performing morphologies, outperforming previous state-of-the-art algorithms. Using significantly less computational budget–approximately 78% fewer simulation steps (20B $\rightarrow$ 4.6B) and 40% fewer training FLOPs per design–Loki explores 780$\sim$ 780$\times$ 780 × more designs than prior evolution-based approaches (4,000 $\rightarrow$ 3M)(Table LABEL:tab:efficiency). Loki discovers varied locomotion strategies including quadrupedal, bipedal, ant-like, crab-like, cheetah, spinner, crawler, and rolling behaviors (Fig. 2). In contrast, prior global competition-based evolutionary methods often converge to a single behavior type–for example, consistently producing only cheetah-like morphologies [4].
Loki results in morphologies that transfer better to new unseen tasks. The final set of evolved designs are trained on a suite of downstream tasks across three domains: agility (bump, patrol, obstacle, exploration), stability (incline), and manipulation (push box incline, manipulation ball). Compared to baseline methods, such as DERL [4] and Map-Elites [24], Loki’s cluster-based co-evolution results in superior adaptability to these challenging unseen tasks (e.g., bump: $981\rightarrow 1908$; push box incline: $1519\rightarrow 3148$; see Fig. 6). Our convergent functions, shared across morphologies, results in high-performing divergent forms that generalize better to new tasks with significantly less compute.
@@ 2 Background and Related Work
Brain-body co-design. The automation of robot behavior is one of the early goals of artificial intelligence [28, 29]. In settings where the robot design is fixed, reinforcement learning (RL) [30, 31, 32, 33] has emerged as a dominant method for optimizing control policies. However, the problem is more complicated when we wish to optimize not only the robot’s policy $\pi\in\Pi$, but also the its physical morphology within some design space $M\in\mathcal{M}$. This design space can be simple real-valued limb length parameters [13], more complex combinatorial structures consisting of link and joint primitives [34], or even the voxelized parameters of soft robots [35]. In these settings, the physical design of a robot and its policy are inherently coupled, as both the transition function $\mathcal{T}\left(s_{t+1}\mid s_{t},a_{t},M\right)$ and the total expected reward $J(\pi,M)=\mathbb{E}_{\pi,M}\sum_{t=0}^{H}\gamma^{t}r_{t}$ are now a function of both action and morphology. This means we must find ways to optimize morphology and behavior jointly: $M^{*}=\text{arg max}_{M}J\left(\pi_{M}^{*},M\right)$ subject to $\pi_{M}^{*}=\text{arg max}_{\pi}{\arg\max}J(\pi,M)$.
@@
Early foundational work optimized morphology and behavior jointly using genetic algorithms [14, 36]. A popular approach is to use bi-level optimization in which an inner RL loop finds the best policies for a given design, and an outer evolutionary loop selects morphologies that produce higher performing behaviors [4]. While the bi-level optimization can be computationally expensive, some authors have proposed solutions such as policy sharing with network structures that can support a variety of morphologies [5, 9]. Others discard the outer evolution loop entirely by incorporating morphology refinement [13] or compositional construction [11, 37] into the RL policy. Between these extremes are approaches that use two reinforcement learning loops that update at different schedules [6]. Recently, researchers have also attempted to produce a diverse set of morphologies rather than a single high-performing design in order to accommodate multiple downstream tasks [38, 4]. Our approach is also in this category and demonstrates substantial improvements over these prior methods.
@@
Optimization and Diversity. Algorithms designed to find multiple diverse solutions to an optimization problem have a long and rich history. Island models [39, 40, 41], fitness sharing [42], tournament selection [43] and niching [44] in genetic algorithms were proposed as different ways to maintain the diversity of a population and prevent premature collapse to a single solution mode. Similar techniques were developed for coevolutionary/multi-agent settings [45, 46].
@@
More recently, Quality Diversity (QD) algorithms [19, 47, 21], attempt to illuminate the search space by discovering a broad repertoire of high-performing yet behaviorally distinct solutions. An important component of algorithm design in this space is deciding how to measure diversity and how to accumulate the resulting collection of solutions. Novelty Search [48, 20] measures diversity as the average distance to nearest neighbors, and adds solutions to an archive if their diversity score is higher than a threshold. MAP-Elites [23, 22] forms a discretized grid of solutions according to a user-defined feature space, and updates this map with the best solution found for each grid cell so far. This grid of solutions has been extended to Voronoi tessellations for high-dimensional feature spaces [49, 50] and combined with evolution strategies [51] and reinforcement learning [52]. QD algorithms can generate a repertoire of diverse solutions, enabling adaptive selection based on changing task demands or environmental conditions [53, 54]. In morphology optimization, evolving a diverse population on a single environment has been shown to yield better generalization and transfer to new tasks [55].
@@ 3 Loki [Uncaptioned image]
We introduce Loki, a compute-efficient co-design framework that discovers diverse, high-performing morphologies (divergent forms) using shared control policies (convergent functions). Leveraging structural commonalities in the design space, we train multi-design policies on clusters of similarly structured morphologies. These shared policies exploit common features, enabling behavior reuse and allowing us to evaluate many more designs without retraining. Morphologies are co-evolved alongside the policies using Dynamic Local Search (DLS) [25, 26, 27], which facilitates broader exploration within each cluster. By localizing the competition [20] within each cluster and having a broader exploration, Loki evolves diverse high-performing solutions that effectively generalize across different unseen tasks. We describe our clustering approach in Sec.3.1, the training of cluster-specific multi-design policies in Sec.3.2, and our co-design framework in Sec. 3.3.
Loki offers three key benefits: Efficient Search. We explore a design space $\sim$780× larger while requiring significantly fewer simulation steps (20B → 4.6B) and less compute per evaluated design (160B → 100B) (Tab. LABEL:tab:efficiency). Robust Evaluation. Single-design policies are narrowly specialized, whereas our multi-design policies are exposed to a broader range of morphologies and behaviors—leading to more robust fitness evaluations grounded in the behavioral landscape of previously encountered agents. Diverse Solutions. Loki independently discovers a range of high-performing designs across clusters, without relying on explicit diversity objectives like novelty search [20].
@@ 3.1 Clustering the Space of Designs
Multi-design policies often struggle to generalize across morphologies with diverse behaviors. To address this, we cluster morphologies that are both structurally and behaviorally similar, and train separate policies for each cluster. Effective generalization within a cluster requires a feature space where distance meaningfully reflects similarity. Each morphology is represented as a sequence of limbs, with a mix of continuous and categorical parameters. Due to the complexity of this representation, clustering in the raw feature space is suboptimal. As shown in Fig. 3, K-means clustering [56] with Euclidean distance over raw parameters groups topologically distinct designs—differing overall structure—into the same cluster, despite similarities in simpler attributes like limb length or radius.
@@
Instead, we train a Transformer-based VAE [57, 58] to encode morphologies into a structured latent space. Each morphology is represented as an $L\times D$ matrix, where $L$ is the maximum number of limbs and $D=47$ is the number of parameters per limb, with both continuous (e.g., limb orientation, length, joint gear) and categorical (e.g., joint type, joint angle, depth) attributes. Limb vectors are ordered using a depth-first search (DFS) traversal [59], starting from the head, which encodes torso type (horizontal/vertical), head density, and radius. If a morphology has fewer than $L$ limbs, the sequence is zero-padded. The transformer encoder [57] maps each morphology to a $(L\times H)$ latent space, where $H<D$. To support decoding, we append an EOS (end-of-sequence) token to mark the final limb and use depth tokens to indicate each limb’s position in the depth-first traversal.
@@
The loss over a batch is defined as $\mathcal{L}^{\text{recon}}=\sum_{i=1}^{L}\mathcal{L}^{\text{recon}}_{i}$, where $\mathcal{L}^{\text{recon}}_{i}=\frac{\sum_{j}\ell_{\text{cont}}(\hat{y}_{ij},y_{ij})\cdot\mathcal{M}^{\text{cont}}_{ij}}{\sum_{j}\mathcal{M}^{\text{cont}}_{ij}}+\frac{\sum_{j}\ell_{\text{cat}}(\hat{y}_{ij},y_{ij})\cdot\mathcal{M}^{\text{cat}}_{ij}}{\sum_{j}\mathcal{M}^{\text{cat}}_{ij}}$ is the reconstruction loss for the $i$-th limb. $\ell_{\text{cont}}$ and $\ell_{\text{cat}}$ are the L2 loss and cross-entropy loss for continuous or categorical attributes. $\mathcal{M}^{\text{cont}}_{ij}$ and $\mathcal{M}^{\text{cat}}_{ij}$ are binary masks showing existence of the $j$-th attribute in the $i$-th limb, and $y_{ij}$ is the ground-truth value. Instead of the standard ELBO objective, we adopt an adaptive scheduling strategy for the KL divergence scaling factor $\beta$ [60]. Our goal is to learn a compact latent representation suitable for clustering. Therefore, we prioritize minimizing the reconstruction loss over enforcing strong regularization towards a standard normal prior. Following  [61], we exponentially decay $\beta$ within a predefined range $[\beta_{\text{min}},\beta_{\text{max}}]$.
@@
We generate 500k unique random morphologies within the UNIMAL space [4] (see Appendix for details). We train a transformer-based VAE (4 layers, 4 heads, latent dimension $H=32$) on these designs using a batch size of 4096, an initial learning rate of $10^{-4}$, and a single A40 GPU for 200 epochs. We then cluster the 500k morphologies into $N_{c}=40$ groups using the K-means algorithm [56] based on Euclidean distance of their corresponding VAE latent codes. A t-SNE [62] visualization of the clusters is shown in Fig. 1-A.
@@
Table 1: Efficiency and Coverage Comparison. Loki explores a much larger design space while requiring significantly fewer simulation steps and less training FLOPs per evaluated design See Appendix for details.

### v2

@@ Abstract
We introduce Loki, a compute-efficient framework for co-designing morphologies and control policies that generalize across unseen tasks. Inspired by biological adaptation—where animals quickly adjust to morphological changes—our method overcomes the inefficiencies of traditional evolutionary and quality-diversity algorithms. We propose learning convergent functions: shared control policies trained across clusters of morphologically similar designs in a learned latent space, drastically reducing the training cost per design. Simultaneously, we promote divergent forms by replacing mutation with dynamic local search, enabling broader exploration and preventing premature convergence. The policy reuse allows us to explore $\sim 780\times$ more designs using 78% fewer simulation steps and 40% less compute per design. Local competition paired with a broader search results in a diverse set of high-performing final morphologies. Using the UNIMAL design space and a flat-terrain locomotion task, Loki discovers a rich variety of designs—ranging from quadrupeds to crabs, bipedals, and spinners—far more diverse than those produced by prior work. These morphologies also transfer better to unseen downstream tasks in agility, stability, and manipulation domains (e.g. $2\times$ higher reward on bump and push box incline tasks). Overall, our approach produces designs that are both diverse and adaptable, with substantially greater sample efficiency than existing co-design methods.
@@ 1 Introduction
Embodied cognition asserts that intelligence emerges not just from the brain but is shaped by the physical embodiment varela2017embodied. This perspective acknowledges that “how a creature behaves is influenced heavily by their body,” which acts as an interface between the controller and the environment stensby2021co. A four legged embodiment gallops while a two legged one simply runs. This adaptation occurs across both evolutionary timescales—as in Darwin’s finches, whose beak shapes evolved to exploit different food sources—and individual lifetimes, through mechanisms like Wolff’s Law, where bones strengthen in response to load wolff2012law. Animals can seamlessly readapt their controllers to accommodate even subtle changes in morphology gupta2021embodied.
Designing—or discovering—optimal embodiments for robots is a difficult combinatorial challenge wang2018nervenet; schaff2019jointly; pathak2019learning; luck2020data; wang2019neural; huang2020one; yuan2021transform2act; kurin2020my; gupta2021embodied; ha2019reinforcement. Co-design, the joint optimization of morphology and control, is typically framed as a bi-level optimization: an outer loop searches for morphologies (often using evolutionary strategies), while an inner loop trains a control policy for each design to evaluate its fitness (performance on a task). Due to the vastness of the design space, exhaustive search is infeasible sims1994evolving; gupta2021embodied. Worse, the fitness landscape is often deceptive goldberg1989genetic, causing evolutionary strategies to converge prematurely to local optima. Escaping such traps requires maintaining diversity across the population hornby2006alps; trujillo2011speciation; trujillo2008discovering. Quality-diversity algorithms address this by promoting exploration of high-performing yet diverse designs pugh2016quality; Lehman2011EvolvingAD; chatzilygeroudis2021quality; Vassiliades2016UsingCV. However, evaluating each candidate is costly, and diversity comes at the expense of the number of designs evaluated in the inner loop Mouret2015IlluminatingSS. Unlike animals—who quickly adapt controllers to new bodies—robots still require retraining from scratch for each morphology, as policies transfer poorly across embodiments gupta2021embodied; mouret2020evolving.
@@
Second, we break away from how most algorithms sample new designs using mutation. Mutating high-performing designs leads to limited exploration. Instead, we sample divergent forms by adopting dynamic local search (DLS) Hoos2015; Hutter2002ScalingAP; Pullan_2006. We maintain a dynamic elite pool per cluster to train a shared policy for the cluster. During training, random designs are sampled from a cluster, evaluated by the current policy, and top ones replace the worst-performing elites in the pool (see Alg. 1). Compared to mutation, which typically yields small, incremental changes, DLS enables broader and more adaptive exploration, reducing the risk of getting stuck in local minima. Since evaluations are made inexpensive with shared policies, we can afford to use random search instead of local mutation, efficiently filtering out low-quality candidates.
We use UNIMAL gupta2021embodied, an expressive design space encompassing approximately $10^{18}$ unique morphologies with fewer than 10 limbs. We evolve morphologies on a simple flat-terrain locomotion task. In nature, locomotion acts as a universal selection pressure across species. It is task-agnostic—helping to prevent overfitting to narrow objectives—and is efficient to simulate and easy to reward.
Beyond being substantially more efficient, Loki discovers both genetically and behaviorally diverse high-performing morphologies, outperforming previous state-of-the-art algorithms. Using significantly less computational budget–approximately 78% fewer simulation steps (20B $\rightarrow$ 4.6B) and 40% fewer training FLOPs per design–Loki explores $\sim 780\times$ more designs than prior evolution-based approaches (4,000 $\rightarrow$ 3M)(Table LABEL:tab:efficiency). Loki discovers varied locomotion strategies including quadrupedal, bipedal, ant-like, crab-like, cheetah, spinner, crawler, and rolling behaviors (Fig. 2). In contrast, prior global competition-based evolutionary methods often converge to a single behavior type–for example, consistently producing only cheetah-like morphologies gupta2021embodied.
Loki results in morphologies that transfer better to new unseen tasks. The final set of evolved designs are trained on a suite of downstream tasks across three domains: agility (bump, patrol, obstacle, exploration), stability (incline), and manipulation (push box incline, manipulation ball). Compared to baseline methods, such as DERL gupta2021embodied and Map-Elites mouret2020evolving, Loki’s cluster-based co-evolution results in superior adaptability to these challenging unseen tasks (e.g., bump: $981\rightarrow 1908$; push box incline: $1519\rightarrow 3148$; see Fig. 6). Our convergent functions, shared across morphologies, results in high-performing divergent forms that generalize better to new tasks with significantly less compute.
@@ 2 Background and Related Work
Brain-body co-design. The automation of robot behavior is one of the early goals of artificial intelligence moravec1980obstacle; center1984shakey. In settings where the robot design is fixed, reinforcement learning (RL) lillicrap2015continuous; Heess2017EmergenceOL; ppo; haarnoja2018soft has emerged as a dominant method for optimizing control policies. However, the problem is more complicated when we wish to optimize not only the robot’s policy $\pi\in\Pi$, but also the its physical morphology within some design space $M\in\mathcal{M}$. This design space can be simple real-valued limb length parameters ha2019reinforcement, more complex combinatorial structures consisting of link and joint primitives dong2023symmetry, or even the voxelized parameters of soft robots li2025generating. In these settings, the physical design of a robot and its policy are inherently coupled, as both the transition function $\mathcal{T}\left(s_{t+1}\mid s_{t},a_{t},M\right)$ and the total expected reward $J(\pi,M)=\mathbb{E}_{\pi,M}\sum_{t=0}^{H}\gamma^{t}r_{t}$ are now a function of both action and morphology. This means we must find ways to optimize morphology and behavior jointly: $M^{*}=\text{arg max}_{M}J\left(\pi_{M}^{*},M\right)$ subject to $\pi_{M}^{*}=\text{arg max}_{\pi}{\arg\max}J(\pi,M)$.
@@
Early foundational work optimized morphology and behavior jointly using genetic algorithms sims1994evolving; sims1994coevolving. A popular approach is to use bi-level optimization in which an inner RL loop finds the best policies for a given design, and an outer evolutionary loop selects morphologies that produce higher performing behaviors gupta2021embodied. While the bi-level optimization can be computationally expensive, some authors have proposed solutions such as policy sharing with network structures that can support a variety of morphologies wang2018nervenet; wang2019neural. Others discard the outer evolution loop entirely by incorporating morphology refinement ha2019reinforcement or compositional construction yuan2021transform2act; lu2025bodygen into the RL policy. Between these extremes are approaches that use two reinforcement learning loops that update at different schedules schaff2019jointly. Recently, researchers have also attempted to produce a diverse set of morphologies rather than a single high-performing design in order to accommodate multiple downstream tasks hejna2021task; gupta2021embodied. Our approach is also in this category and demonstrates substantial improvements over these prior methods.
@@
Optimization and Diversity. Algorithms designed to find multiple diverse solutions to an optimization problem have a long and rich history. Island models tanese1989distributed; belding1995distributed; whitley1997island, fitness sharing goldberg1987genetic, tournament selection oei1991tournament and niching sareni1998fitness in genetic algorithms were proposed as different ways to maintain the diversity of a population and prevent premature collapse to a single solution mode. Similar techniques were developed for coevolutionary/multi-agent settings cartlidge2003caring; mitchell2006role.
@@
More recently, Quality Diversity (QD) algorithms pugh2016quality; Cully2017QualityAD; chatzilygeroudis2021quality, attempt to illuminate the search space by discovering a broad repertoire of high-performing yet behaviorally distinct solutions. An important component of algorithm design in this space is deciding how to measure diversity and how to accumulate the resulting collection of solutions. Novelty Search krvcah2010solving; Lehman2011EvolvingAD measures diversity as the average distance to nearest neighbors, and adds solutions to an archive if their diversity score is higher than a threshold. MAP-Elites Mouret2015IlluminatingSS; Vassiliades2016UsingCV forms a discretized grid of solutions according to a user-defined feature space, and updates this map with the best solution found for each grid cell so far. This grid of solutions has been extended to Voronoi tessellations for high-dimensional feature spaces vassiliades2017using; vassiliades2018discovering and combined with evolution strategies colas2020scaling and reinforcement learning pierrot2023evolving. QD algorithms can generate a repertoire of diverse solutions, enabling adaptive selection based on changing task demands or environmental conditions gaier2017data; cully2015robots. In morphology optimization, evolving a diverse population on a single environment has been shown to yield better generalization and transfer to new tasks nordmoen2021map.
@@ 3 Loki [Uncaptioned image]
We introduce Loki, a compute-efficient co-design framework that discovers diverse, high-performing morphologies (divergent forms) using shared control policies (convergent functions). Leveraging structural commonalities in the design space, we train multi-design policies on clusters of similarly structured morphologies. These shared policies exploit common features, enabling behavior reuse and allowing us to evaluate many more designs without retraining. Morphologies are co-evolved alongside the policies using Dynamic Local Search (DLS) Hoos2015; Hutter2002ScalingAP; Pullan_2006, which facilitates broader exploration within each cluster. By localizing the competition Lehman2011EvolvingAD within each cluster and having a broader exploration, Loki evolves diverse high-performing solutions that effectively generalize across different unseen tasks. We describe our clustering approach in Sec.3.1, the training of cluster-specific multi-design policies in Sec.3.2, and our co-design framework in Sec. 3.3.
Loki offers three key benefits: Efficient Search. We explore a design space $\sim$780× larger while requiring significantly fewer simulation steps (20B → 4.6B) and less compute per evaluated design (160B → 100B) (Tab. LABEL:tab:efficiency). Robust Evaluation. Single-design policies are narrowly specialized, whereas our multi-design policies are exposed to a broader range of morphologies and behaviors—leading to more robust fitness evaluations grounded in the behavioral landscape of previously encountered agents. Diverse Solutions. Loki independently discovers a range of high-performing designs across clusters, without relying on explicit diversity objectives like novelty search Lehman2011EvolvingAD.
@@ 3.1 Clustering the Space of Designs
Multi-design policies often struggle to generalize across morphologies with diverse behaviors. To address this, we cluster morphologies that are both structurally and behaviorally similar, and train separate policies for each cluster. Effective generalization within a cluster requires a feature space where distance meaningfully reflects similarity. Each morphology is represented as a sequence of limbs, with a mix of continuous and categorical parameters. Due to the complexity of this representation, clustering in the raw feature space is suboptimal. As shown in Fig. 3, K-means clustering lloyd1982kmeans with Euclidean distance over raw parameters groups topologically distinct designs—differing overall structure—into the same cluster, despite similarities in simpler attributes like limb length or radius.
@@
Instead, we train a Transformer-based VAE transformers; vae to encode morphologies into a structured latent space. Each morphology is represented as an $L\times D$ matrix, where $L$ is the maximum number of limbs and $D=47$ is the number of parameters per limb, with both continuous (e.g., limb orientation, length, joint gear) and categorical (e.g., joint type, joint angle, depth) attributes. Limb vectors are ordered using a depth-first search (DFS) traversal cormen2022introduction, starting from the head, which encodes torso type (horizontal/vertical), head density, and radius. If a morphology has fewer than $L$ limbs, the sequence is zero-padded. The transformer encoder transformers maps each morphology to a $(L\times H)$ latent space, where $H<D$. To support decoding, we append an EOS (end-of-sequence) token to mark the final limb and use depth tokens to indicate each limb’s position in the depth-first traversal.
@@
The loss over a batch is defined as $\mathcal{L}^{\text{recon}}=\sum_{i=1}^{L}\mathcal{L}^{\text{recon}}_{i}$, where $\mathcal{L}^{\text{recon}}_{i}=\frac{\sum_{j}\ell_{\text{cont}}(\hat{y}_{ij},y_{ij})\cdot\mathcal{M}^{\text{cont}}_{ij}}{\sum_{j}\mathcal{M}^{\text{cont}}_{ij}}+\frac{\sum_{j}\ell_{\text{cat}}(\hat{y}_{ij},y_{ij})\cdot\mathcal{M}^{\text{cat}}_{ij}}{\sum_{j}\mathcal{M}^{\text{cat}}_{ij}}$ is the reconstruction loss for the $i$-th limb. $\ell_{\text{cont}}$ and $\ell_{\text{cat}}$ are the L2 loss and cross-entropy loss for continuous or categorical attributes. $\mathcal{M}^{\text{cont}}_{ij}$ and $\mathcal{M}^{\text{cat}}_{ij}$ are binary masks showing existence of the $j$-th attribute in the $i$-th limb, and $y_{ij}$ is the ground-truth value. Instead of the standard ELBO objective, we adopt an adaptive scheduling strategy for the KL divergence scaling factor $\beta$ higgins2017betavae. Our goal is to learn a compact latent representation suitable for clustering. Therefore, we prioritize minimizing the reconstruction loss over enforcing strong regularization towards a standard normal prior. Following  zhang2024mixedtype, we exponentially decay $\beta$ within a predefined range $\beta_{\text{min}},\beta_{\text{max}}]$.
@@
We generate 500k unique random morphologies within the UNIMAL space gupta2021embodied (see Appendix C for details). We train a transformer-based VAE (4 layers, 4 heads, latent dimension $H=32$) on these designs using a batch size of 4096, an initial learning rate of $10^{-4}$, and a single A40 GPU for 200 epochs. We then cluster the 500k morphologies into $N_{c}=40$ groups using the K-means algorithm lloyd1982kmeans based on Euclidean distance of their corresponding VAE latent codes. A t-SNE van2008visualizing visualization of the clusters is shown in Fig. 1-A.
@@
Table 1: Efficiency and Coverage Comparison. Loki explores a much larger design space while requiring significantly fewer simulation steps and less training FLOPs per evaluated design See Appendix F for details.